OpenStudy (anonymous):

hi, how do you geometrically prove that two vectors are parallel without using the dot product?

OpenStudy (anonymous):

when there coefficient are in proportional

OpenStudy (anonymous):

sweet, thanks!

OpenStudy (anonymous):

consider , two vectors u =ai+bj+ck v =xi+yj+zk then , u and v are parallel when a/x=b/y=c/z